# Exploiting the Layered Intrinsic Dimensionality of Deep Models for Practical Adversarial Training

### v1

@@ Abstract
Despite being a heavily researched topic, Adversarial Training (AT) is rarely, if ever, deployed in practical AI systems for two primary reasons: (i) the gained robustness is frequently accompanied by a drop in generalization and (ii) generating adversarial examples (AEs) is computationally prohibitively expensive. To address these limitations, we propose SMAAT  a new AT algorithm that leverages the manifold conjecture [49], stating that off-manifold AEs lead to better robustness while on-manifold AEs result in better generalization. Specifically, SMAAT aims at generating a higher proportion of off-manifold AEs by perturbing the intermediate deepnet layer with the lowest intrinsic dimension. This systematically results in better scalability compared to classical AT as it reduces the PGD chains length required for generating the AEs. Additionally, our study provides, to the best of our knowledge, the first explanation for the difference in the generalization and robustness trends between vision and language models, i.e., AT results in a drop in generalization in vision models whereas, in encoder-based language models, generalization either improves or remains unchanged. We show that vision transformers and decoder-based models tend to have low intrinsic dimensionality in the earlier layers of the network (more off-manifold AEs), while encoder-based models have low intrinsic dimensionality in the later layers. We demonstrate the efficacy of SMAAT  on several tasks, including robustifying (i) sentiment classifiers, (ii) safety filters in decoder-based models, and (iii) retrievers in RAG setups. SMAAT requires only 25-33% of the GPU time compared to standard AT, while significantly improving robustness across all applications and maintaining comparable generalization.
@@ 1 Introduction
Adversarial training (AT) has been the most effective approach for training deep neural network models to be robust to small input perturbations [6, 28]. It is formulated as a min-max optimization problem where the outer minimization is over the parameter space and the inner-maximization searches for the worst-case input space perturbation. For deepnets, the inner maximization is classically solved approximately using several iterations of the projected gradient descent method (PGD, Madry et al. [38]). However, AT suffers from at least two major limitations. First, the impact of improving robustness on generalization is not well understood [55]. For instance, in vision models, AT leads to a drop in generalization (e.g., 8% drop on CIFAR10 [48]), whereas in encoder-based language models (enc-LLMs) [^1] We will use ”enc-LLMs” to refer to encoder-based language models such as BERT and RoBERTa, and ”dec-LLMs” for decoder-based language models such as LLaMA, it either improves generalization or remains unchanged (e.g., 1% improvement on AGNEWS [80]). Moreover, robustness improvement is consistently more significant in vision (e.g., 40% on CIFAR10) than in enc-LLMs (e.g., 10% on AGNEWS). While there has been a lot of effort to understand the drop in generalization in vision models [38, 65, 3, 73, 13], there has not been any work understanding the inherent difference in the robustness-generalization relationship in vision, dec-LLMs and enc-LLMs. Second, AT is associated with a high computational overhead as it requires several forward and backward passes through the network. This significantly limits AT scalability [46] to increasingly wider and deeper models [15, 8]. To address this, recent research efforts focused on minimizing the number of PGD steps, often through reusing or accumulating gradients during parameter updates [48, 72, 80]. Yet, these methods still require full-length network passes and, despite improvement, continue to be computationally demanding.
(a) Classification
(b) RAG
(c) Safety Filter
@@
Figure 1: Comparison of SMAAT robustness, generalization, and run time (marker size) against baselines for robustifying (a) topic classifiers, (b) retriever models in the Retrieval Augmented Generation (RAG) setup and (c) safety filters for decoder-based LLMs. SMAAT significantly enhances model robustness compared to seven different baselines, while maintaining nearly the same clean accuracy. Additionally, it greatly improves the scalability of adversarial training.
@@
In this work, we first identify a root cause of the difference in the robustness-generalization trends in foundational models. We show that the intrinsic dimension (ID) of the data manifold in the first layer is much higher in enc-LLMs compared to vision models/dec-LLM (see Figure 2) which results in AEs being more on-manifold (ONM-AEs) in enc-LLMs and more off-manifold (OFM-AEs) in vision models and dec-LLMs. In accordance with the manifold conjecture [18, 50, 21], we find that OFM-AEs lead to better robustness while more ONM-AEs lead to better generalization. To the best of our knowledge, this is the first explanation for the difference in the robustness magnitudes across vision models and enc-LLMs. Our findings are also consistent with YOPO [72] and TMD [40]. Specifically, YOPO highlights the critical role of the first layer in vision models for AT, while TMD uses the last layer to detect AEs in enc-LLMs.
@@
To address the second limitation, i.e., scalability, we leverage our aforementioned insights on the impact of the manifold ID on robustness and generalization and hypothesize that perturbing the intermediate layer $l$ with the highest off-manifold AE ratio (equivalently lowest ID) should achieve high robustness at low computational cost. Intuitively, a lower ID corresponds to a higher proportion of OFM-AEs generation and a greater improvement in robustness by the manifold conjecture. Additionally, layers closer to the output result in shorter PGD chains, leading to efficient AT. To this end, we propose SMAAT, a Scalable Manifold Aware AT approach that applies AT at the layer with the highest proportion of OFM-AEs. We found that this critical layer is consistently the last layer across enc-LLMs on several applications. Hence, for enc-LLMs, SMAAT leads to a significant speed-up of AT by avoiding a full backward pass as it calculates the gradients only until the last layer rather than the entire network. Also, empirically, we show that we achieve a higher robustness. Yet, this ID trend is different for vision or dec-LLMs, where we find that the first layer is always the one with the lowest ID. So, SMAAT in this case would be equivalent to standard adversarial training. Nevertheless, we argue that there is significant value in robustifying enc-LLMs as they are increasingly being used to augment the capabilities (e.g., retrievers in RAG setups [23]) or improve the safety (e.g., safety filters [27]) in dec-LLMs.
Figure 2: Overview. First, we study the differences in robustness and generalization trends in AT in encoder-based (enc-), decoder-based (dec-) large language models (LLMs), and vision models (Left). We observe that the initial layer intrinsic dimension is significantly lower in dec-LLMs and vision leading to a higher proportion of off-maifold AEs and consequently better robustness unlike to enc-LLMs. Motivated by this insight, we incorporate this in a new manifold-aware AT paradigm, SMAAT. SMAAT proceeds in two phases. First, it searches for the layer $l^{*}$ with the lowest ID. Second, SMAAT trains the network adversarially by generating samples from $l^{*}$ as opposed to $l_{1}$, as classically. This results in, $\mathcal{O}\left(P\times(n-l^{*})\right)$ gain in run-time complexity per training iteration (Right)
@@
We evaluated the effectiveness of SMAAT on robustifying (1) classifiers, (2) safety filters, and (3) retrievers in Retrieval-Augmented Generation (RAG). SMAAT achieved state-of-the-art (SOTA) runtime and robustness results on all tasks, while maintaining clean accuracy (generalization) comparable to standard training (Fig.1). Specifically, in sentiment and content classification tasks, SMAAT improved the robustness of enc-LLMs on AGNEWS [76], IMDB [36], and YELP [74] datasets by 8.6%, 15.7%, and 28.8% for BERT [15] and by 6.0%, 5.8%, and 19.0% for RoBERTa [34], respectively. For safety filtering, SMAAT accurately identified harmful prompts generated by the GCG attack [81], achieving 97-100% accuracy. In RAG experiments, SMAAT significantly enhanced the robustness of the Contrevier model [23] on the RAG setup, achieving over 80% robustness against poisoning attacks [78]. Furthermore, SMAAT required only about 25-33% of the GPU time compared to the standard AT.
@@
    SMAAT, a novel adversarial training algorithm that leverages the intrinsic dimensionality across layers of foundational models to control robustness and generalization.
@@
    An explanation for the discrepancy in the robustness and generalization trends in foundational models.
@@
    Comprehensive experiments demonstrating enhancements in robustness and scalability, while maintaining accuracy in classification and retrieval tasks compared to standard adversarial training.
@@ 2 Related Work
Adversarial Training (AT) aims at robustifying a model against AEs which are imperceptibly perturbed inputs that can lead to incorrect predictions. Formally, AT is framed as a robust optimization program. Specifically, it seeks the optimal parameters $\theta^{*}$ of a classifier $f_{\theta}(x)$ that are robust to perturbations $\delta$ within a norm ball, i.e.,
@@
where $\ell$ is a loss function (e.g., cross-entropy loss) and $\mathcal{D}=\{x,y\}_{i=1}^{|\mathcal{D}|}$ is the training data. It has been shown that the saddle-point problem can be solved reliably with SGD for the outer minimization and with project-gradient descent (PGD) for the inner maximization [38]. PGD starts at randomly initialized perturbations in the $\epsilon$-ball and iteratively applies a gradient ascent update followed by a projecting onto the $\epsilon-$ball $P$ times. $P$-step PGD AT is an order of magnitude more expensive than standard training for deep networks as the $P$-step PGD requires $P$ forward-backward passes through the network, while the standard SGD update requires only one.
@@
Manifold-Based Defenses. The manifold conjecture stands as one of the most compelling explanations for the susceptibility of deep neural networks to adversarial samples [57, 21, 50]. The conjecture posits that data resides on a low-dimensional manifold within a high-dimensional representation space and that the deepnet model learns to approximate this manifold. Consequently, an off-manifold sample, deviating from this foundational manifold, leads to undefined behavior of the model. This conjecture has inspired a novel line of defenses against adversarial attacks on images [45, 39, 53, 47] and text [40]. These methods approximate the data manifold and, during testing, project samples onto this manifold to either detect or correctly classify. Differently, we leverage the manifold conjecture during training to improve both robustness and scalability.
@@
Robustness and Generalization Trends in AT. AT is recognized for enhancing model robustness in both vision and enc-LLMs [73, 4]. While this improvement in robustness comes at the cost of increased generalization error in vision models [73], AT enhances generalization in enc-LLMs [4, 20]. There has been recently efforts relating robustness to OFM-AEs and generalization to ONM-AEs [56, 70]. We extend this line of research by connecting these trends to the intrinsic dimensionality of intermediate layers of a network.
@@
Scalable AT. Different optimizations have been proposed to mitigate the cost of the PGD additional $P$ forward-backward passes for AEs generation, including (i) replacing multi-step PGD with a single-step FGSM [48]; (ii) omitting redundant computations during PGD [72]; (iii) combining FGSM AT with random initialization [68]. While these approaches alleviate the PGD overhead, they also come with limitations. As a solution, FreeLB [80] proposes accumulating model parameter gradients over multiple batches. In our work, we achieve scalability through an alternative method, i.e., by leveraging the manifold conjecture to perturb intermediate layers and hence reduce the length of backward-forward chains in PGD.
ID Estimation. ID corresponds to the minimum number of variables needed in a maximum representation of data. Various techniques have been developed to estimate ID, including (i) Two-Nearest Neighbors (twoNN), which examines distances to the two nearest neighbors [19], (ii) Singular Value Decomposition (SVD), used for projection error estimation by decomposing a matrix into singular values [54], (iii) Maximum Likelihood Estimation (MLE), which maximizes the likelihood of observed distances [43], and (iv) Correlation Dimension, which estimates ID based on the scaling behavior of pairwise distances [22]. In our work, we utilize twoNN for ID estimation due to its robustness to non-linearity of the data manifold. Specifically, it analyzes the ratio $\mu$ of distances from each point to its two nearest neighbors. The empirical distribution function of $\mu$ is modeled using a Pareto distribution model, expressed as $f(\mu|I)=I\mu^{-(I+1)}$, where $I$ is the ID of the manifold [19]. However, since twoNN is not applicable for estimating the OFM/ONM proportion of AEs, we utilize the SVD projection error estimation approach for this purpose.

### v2

# Explaining the role of Intrinsic Dimensionality in Adversarial Training

@@ Abstract
Adversarial Training (AT) impacts different architectures in distinct ways: vision models gain robustness but face reduced generalization, encoder-based models exhibit limited robustness improvements with minimal generalization loss, and recent work in latent-space adversarial training (LAT) demonstrates that decoder-based models achieve improved robustness by applying AT across multiple layers. We provide the first explanation for these trends by leveraging the manifold conjecture: off-manifold adversarial examples (AEs) enhance robustness, while on-manifold AEs improve generalization. We show that vision and decoder-based models exhibit low intrinsic dimensionality in earlier layers (favoring off-manifold AEs), whereas encoder-based models do so in later layers (favoring on-manifold AEs). Exploiting this property, we introduce SMAAT, which improves the scalability of AT for encoder-based models by perturbing the layer with the lowest intrinsic dimensionality. This reduces the projected gradient descent (PGD) chain length required for AE generation, cutting GPU time by 25–33% while significantly boosting robustness. We validate SMAAT across multiple tasks, including text generation, sentiment classification, safety filtering, and retrieval augmented generation setups, demonstrating superior robustness with comparable generalization to standard training.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Impact of applying LAT at different layers of the LLaMA-2 model, illustrating the relationship between Intrinsic Dimensionality (background color), Generalization (blue), Robustness (orange), and Off-Manifold Ratio (green, based on reconstruction error). Markers show the average of measured values across multiple training configurations; lines depict overall trends. The off-manifold ratio measures the percentage of adversarial examples that fall outside the data manifold using reconstruction error. As we move to deeper layers, the Intrinsic Dimensionality increases, resulting in a decrease in the off-manifold ratio. According to the manifold conjecture, this leads to an increase in generalization (more on-manifold samples) and a decrease in robustness.
Adversarial Training (AT) is the most effective approach for improving the robustness of deep neural networks against small input perturbations (Bai et al., 2021; Kurakin et al., 2017). It is formulated as a min-max optimization problem, where the outer minimization optimizes model parameters, and the inner maximization seeks worst-case input perturbations. In deep networks, the inner maximization is typically solved approximately using multiple iterations of projected gradient descent (PGD, Madry et al. (2017)). However, the trade-off between robustness and generalization remains poorly understood. For instance, AT reduces generalization in vision models, leading to an 8% drop on CIFAR-10 (Shafahi et al., 2019), whereas encoder-based language models often retain or even improve generalization, achieving a 1% gain on AGNEWS (Zhu et al., 2020). Moreover, robustness gains are significantly higher in vision models (e.g., 40% on CIFAR-10) compared to encoder-based models (e.g., 10% on AGNEWS). Recent work on Latent Adversarial Training (LAT) suggests that applying AT across multiple layers yields better robustness and generalization than focusing on a single layer (Sheshadri et al., 2024). While prior research has investigated generalization loss in vision models (Madry et al., 2017; Wang et al., 2019b; Altinisik et al., 2023a; Zhang et al., 2019b; Cheng et al., 2022), no study has systematically explored the robustness-generalization dynamics across vision models (CNNs and Vision Transformers), decoder-based (dec-LLMs), and encoder-based (enc-LLMs) language models. Additionally, the high computational cost of AT limits its practical deployment. Recent efforts to reduce the number of PGD steps—by reusing or accumulating gradients during updates (Shafahi et al., 2019; Zhang et al., 2019a; Zhu et al., 2020)—have improved efficiency but still require full network passes, making AT computationally expensive.
@@
Figure 2: Comparison of SMAAT robustness (x-axis), generalization (y-axis), and run time (marker size) against baselines for robustifying (a) topic classifiers, (b) retriever models in the Retrieval Augmented Generation (RAG) setup and (c) safety filters for decoder-based LLMs. SMAAT significantly enhances model robustness compared to seven different baselines, while maintaining nearly the same clean accuracy. Besides, it is significantly more scalable than AT (marker size).
@@
In this work, we investigate how differences in robustness and generalization trends across foundational models relate to the intrinsic dimensionality (ID) of the data manifold. The data manifold is a potentially non-linear subspace spanned by the dataset, and its dimensionality influences whether adversarial examples (AEs) lie on the manifold (on-manifold AEs) or fall outside it (off-manifold AEs) during training. We show that the ID of the data manifold in the first layer is much higher in enc-LLMs compared to vision models and dec-LLMs which results in AEs being more on-manifold (ONM-AEs) in enc-LLMs and more off-manifold (OFM-AEs) in vision models and dec-LLMs. In accordance with the manifold conjecture (Ethayarajh, 2019; Shamir et al., 2021; Gilmer et al., 2018), we find that OFM-AEs lead to better robustness, while more ONM-AEs lead to better generalization (Fig. 1, with details provided in Section 4). To the best of our knowledge, this is the first explanation for the difference in the robustness magnitudes across vision models and enc-LLMs. Our findings are also consistent with YOPO (Zhang et al., 2019a) and TMD (Minh & Luu, 2022). Specifically, YOPO highlights the critical role of the first layer in vision models for AT, while TMD uses the last layer to detect AEs in enc-LLMs.
@@
We further leverage our insights on the impact of the manifold ID on robustness and generalization and hypothesize that perturbing the intermediate layer $l$ with the highest off-manifold AE ratio (equivalently lowest ID) should achieve high robustness at low computational cost. Intuitively, a lower ID corresponds to a higher proportion of OFM-AEs generation and a greater improvement in robustness by the manifold conjecture. Additionally, layers closer to the output result in shorter PGD chains, leading to efficient AT. To this end, we propose SMAAT [^1] The code is publicly available at: https://github.com/EnesAltinisik/SMAAT-25/tree/main, a Scalable Manifold Aware AT approach that applies AT at the layer with the highest proportion of OFM-AEs. We found that this critical layer is consistently the last layer across enc-LLMs on several applications. Hence, for enc-LLMs, SMAAT leads to a significant speed-up of AT by avoiding a full backward pass as it calculates the gradients only until the last layer rather than the entire network. Also, empirically, we show that we achieve a higher robustness. Yet, this ID trend is different for vision or dec-LLMs, where we find that the first layer is always the one with the lowest ID. In these cases, SMAAT effectively reduces to standard AT.
Figure 3: Left: In classical AT, adversarial examples (AEs) are created in the data layer. For encoder LLMs, the intrinsic dimensionality tends to be high in the initial layers and therefore the AEs created tend to be on-manifold which results in better generalization. In Vision and decoder LLMs, we observe the opposite behavior and AEs tend to be off-manifold resulting in better robustness. Right: The key idea of SMAAT is to create AEs in intermediate layers where the intrinsic dimensionality is low and AEs will tend to be off-manifold. This results in better robustness while (surprisingly) maintaining generalization. The speed-up in SMAAT is due to the fact we need shorter backprop chains to create AEs in intermediate layers.
@@
Given enc-LLMs continue to play a crucial role in machine learning pipelines, there is significant value in effectively enhancing their robustness. To this end, we evaluated SMAAT in improving the robustness of (1) classifiers, (2) safety filters, and (3) retrievers within Retrieval-Augmented Generation (RAG). SMAAT achieved state-of-the-art (SOTA) runtime and robustness results on all tasks, while maintaining clean accuracy (generalization) comparable to standard training. Specifically, in sentiment and content classification tasks, SMAAT improved the robustness of enc-LLMs on AGNEWS (Zhang et al., 2015), IMDB (Maas et al., 2011), and YELP (Zhang et al., 2015) datasets by 8.6%, 15.7%, and 28.8% for BERT (Devlin et al., 2019) and by 6.0%, 5.8%, and 19.0% for RoBERTa (Liu et al., 2019), respectively. For safety filtering, SMAAT accurately identified harmful prompts generated by GCG (Zou et al., 2023), achieving 97-100% accuracy. In RAG experiments, SMAAT significantly enhanced the robustness of the Contrevier model (Izacard et al., 2021) on the RAG setup, achieving over 80% robustness against poisoning attacks (Zhong et al., 2023). Besides, SMAAT required only about 25-33% of the GPU time compared to the standard AT. A summary of all results is presented in Fig. 2.
@@
    An explanation for the discrepancy in the robustness and generalization trends in foundation models.
@@
    SMAAT, a novel AT algorithm that leverages the intrinsic dimensionality across layers of foundational models to control robustness, generalization and scalability.
@@
    Comprehensive experiments demonstrating enhancements in robustness and scalability while keeping accuracy in classification and retrieval tasks compared to standard AT.
@@ 2 Related Work
Adversarial Training (AT) aims at robustifying a model against AEs which are imperceptibly perturbed inputs that can lead to incorrect predictions. Formally, AT seeks optimal parameters $\theta^{*}$ for a classifier $f_{\theta}(x)$ that remains robust to perturbations $\delta$ within a norm ball:
@@
where $\ell$ is the loss function and $\mathcal{D}=\{(x,y)\}_{i=1}^{|\mathcal{D}|}$ represents the training data. The outer minimization is typically solved using stochastic gradient descent (SGD), while the inner maximization is addressed via projected gradient descent (PGD) (Madry et al., 2017). LAT extends AT by applying adversarial perturbations to the model’s latent representations instead of its inputs (Casper et al., 2024; Sheshadri et al., 2024). Given a model with parameters $\theta=(\theta_{1},\theta_{2})$, it computes $h_{\theta_{2}}\circ g_{\theta_{1}}$, where $g_{\theta_{1}}$ is a feature extractor that maps inputs to latent representations $\ell_{i}=g_{\theta_{1}}(x_{i})$, and $h_{\theta_{2}}$ maps these latents to predictions $\hat{y}_{i}=h_{\theta_{2}}(\ell_{i})$. The standard AT objective under an $L_{p}$-norm constraint of $\epsilon$ is:
$\min_{\theta}\sum_{i}\max_{\delta_{i}}\;\mathcal{L}(h_{\theta_{2}}(g_{\theta_{1}}(x_{i}+\delta_{i})),y_{i})$
Similar to input-space AT, the inner maximization finds the worst-case perturbation $\delta_{i}$, while the outer minimization updates $\theta$, both solved via gradient-based methods. However, AT with $P$-step PGD is significantly more computationally expensive than standard training, as it requires $P$ forward-backward passes per update, compared to just one in standard SGD.
Adversarial Attacks on LLMs. Adversarial attacks on LLMs are more challenging than on vision models due to the discrete nature of text and tokenization. Early attacks on enc-LLMs used word substitutions guided by embedding similarity (Jin et al., 2020b), synonymity (Zang et al., 2019), or masked language models (Li et al., 2020b), primarily to flip classifier outputs. In contrast, recent attacks on dec-LLMs focus on alignment breaking objectives by appending adversarial prefixes or suffixes. For example, AutoDAN (Liu et al., 2023) and PAIR (Wei et al., 2023) craft prompts that encourage harmful completions, while Greedy Coordinate Gradient (GCG) attack (Zou et al., 2023) appends a gradient based generated adversarial suffix to user inputs. GCG combines affirmative prompting (Wei et al., 2023; Carlini et al., 2023) with greedy and gradient-based discrete optimization (Shin et al., 2020), and is notable for its strong transferability across prompts and models.
@@
Manifold-Based Defenses. The manifold conjecture stands as one of the most compelling explanations for the susceptibility of deep neural networks to AEs (Tanay & Griffin, 2016; Gilmer et al., 2018; Shamir et al., 2021). The conjecture posits that data resides on a low-dimensional manifold within a high-dimensional representation space and that the deepnet learns to approximate this manifold. Consequently, an off-manifold sample, deviating from this foundational manifold, leads to undefined behavior of the model. This conjecture has inspired a novel line of defenses against adversarial attacks on images (Samangouei et al., 2018; Meng & Chen, 2017; Song et al., 2017; Schott et al., 2018) and text (Minh & Luu, 2022). These methods approximate the data manifold and, during testing, project samples onto this manifold to either detect or correctly classify AEs. Differently, we leverage the manifold conjecture during training to improve both robustness and scalability of AT.
@@
Robustness and Generalization Trends in AT. AT is recognized for enhancing model robustness in both vision and enc-LLMs (Zhang et al., 2019b; Altinisik et al., 2023b). While this improvement in robustness comes at the cost of increased generalization error in vision models (Zhang et al., 2019b), AT enhances generalization in enc-LLMs (Altinisik et al., 2023b; Gan et al., 2020). There has been recently efforts relating robustness to OFM-AEs and generalization to ONM-AEs (Stutz et al., 2019b; Xiao et al., 2022). We extend this line of research by connecting these trends to the intrinsic dimensionality of intermediate layers of a network.
Figure 4: The ID (row 1) trend follows the inverse OFM-/ONM-AEs ratio (row 2) trend. The average projection error ($e_{l}^{k}$) is used as a proxy for estimating the OFM-/ONM- AEs ratio. The ID is computed using the twoNN approach. Enc-LLMs (BERT, RoBERTa) have decreasing ID and OFM-AEs proportions trends unlike vision models and dec-LLMs that have increasing ID and ONM-AEs trends.
@@
Scalable AT. Different optimizations have been proposed to mitigate the cost of the PGD additional forward-backward passes, including (i) replacing multi-step PGD with a single-step FGSM (Shafahi et al., 2019); (ii) omitting redundant computations during PGD (Zhang et al., 2019a); (iii) combining FGSM with random initialization (Wong et al., 2020). While these approaches alleviate the PGD overhead, they also come with limitations. As a solution, FreeLB (Zhu et al., 2020) proposes accumulating model parameter gradients over multiple batches. We achieve scalability through an alternative method, i.e., by leveraging the manifold conjecture to perturb intermediate layers and thus reduce the length of backward-forward PGD chains.
ID Estimation. It is well known that most real world data lives in a low-dimensional space relative to the ambient space where it is defined. The ID of data is the minimum number of variables necessary to characterize important properties of the data. Singular Value Decomposition (SVD) is a well known method to estimate the ID assuming the data lives on a linear manifold (Stewart, 1993). While there have been many proposals to estimate the ID in non-linear settings, we will use the relatively recent twoNN ID-estimator based on the observation that for every data point $x$, the ratio $\mu$ of distance of $x$ to it’s second and first nearest neighbor follows a Pareto distribution $f(\mu|I)=I\mu^{-(I+1)}$, where $I$ is the ID. It can be estimated as $I=\frac{\log(1-F(\mu))}{\log(\mu)}$, where $F(\mu)$ is the empirical CDF (Facco et al., 2017).